OpenStudy (anonymous):

What is the quotient of the complex numbers below? 3+2i/1-5i A.) -7/26 + 17/26i B.) 13-13i C.)3-2/5i D.) -24 -10i

OpenStudy (anonymous):

In case you didn't understand.

OpenStudy (igreen):

\(\dfrac{3+2i}{1-5i} = -\dfrac{7}{26}+\dfrac{17}{26}i\)

OpenStudy (anonymous):

-7/26 + 17/26i Would you like to see why?

OpenStudy (anonymous):

Thanks & yes please